OpenStudy (anonymous):

4/6-x+x^2/6-x=2 solve for x and show work

OpenStudy (bibby):

does that say\(\large \frac{4}{6-x}+\frac{x^2}{6-x}=2\) ?

OpenStudy (anonymous):

yes!

OpenStudy (bibby):

\(\large \frac{a}{b}+\frac{c}{b}=\frac{a+c}{b}\)

OpenStudy (anonymous):

huh

OpenStudy (bibby):

if 2 fractions have the same denominator and are being added or subtracted, you can just put t hem together over the same denominator. \(\huge \frac{6}{x+2}-\frac{7}{x+2}=\frac{6-7}{x+2}\)